Please help with the chart at least.

On January 1, 20X1. Par Company purchased all the outstanding stock of South Bay Company, located in Canada, for $129,600. Onb. Par uses the fully adjusted equity method to account for its investment. Provide the entries that it would record in 20x1Record the parents share (100%) of the translation adjustment from the translation of the subsidiarys accounts on December
Answer #1

Part A

Canadian Dollars

Exchange rate

U.S. Dollars

Investment cost

C

160000

0.81

$129600

Book value of investment on January 1, 20X1

81000

0.81

65610

Differential

C

79000

63990

129600/0.81 = 160000

Canadian Dollars

Exchange rate

U.S. Dollars

Plant and equipment

Trademark

Plant and equipment

Trademark

Income statement:

Differential at date of acquisition:

10700

68300

0.81

8667

55323

Amortization this period (10 years)

(1070)

(6830)

0.76

(813)

(5191)

Remaining balance:

9630

61470

7854

50132

Balance sheet

Remaining balance on 12/31/x1 translated at year-end exchange rates:

9630

61470

0.64

6163

39341

Difference to OCI – translation adjustment:

1691

10791

Differential = investment cost – book value of investment = (129600/0.81) – 81000 = 160000-81000 = $79000

Trademark (Differential) = 79000-10700 = 58300

Part B

No.

General journal

Debit

Credit

A

Investment in south bay company

129600

Cash

129600

(to record the acquisition of South Bay company)

B

Investment in south bay company

11400

Income from subsidiary (15000*0.76)

11400

(to record the equity in income of the subsidiary)

C

Foreign currency units (C$)

5475

Investment in south bay company (7300*0.75)

5475

(to record the dividend from the foreign subsidiary)

D

Income from subsidiary (813+5191)

6004

Investment in south bay company

6004

(to record the amortization of differential)

E

Other comprehensive income – translation adjustment (1691+10791)

12482

Investment in south bay company

12482

(to record entry to recognize the translation adjustment on the differential)

Part C

PAR COMPANY AND SUBSIDIARY

Proof of Translation Adjustment

Year ended December 31, 20X1

Canadian Dollars

Exchange rate

U.S. Dollars

Net assets at begginnig of year, 1/1/X1

81000

0.81

65610

Adjustments for changes in assets portion during year:

Net income for year

15000

0.76

11400

Dividends paid

(7300)

0.75

(5475)

Net assets transferred at rates in effect for those items

71535

Net assets at end of year

88700

0.64

56768

Change in other comprehensive income – translation adjustment during year – net decrease (debit)

14767

No.

General journal

Debit

Credit

A

Other comprehensive income – translation adjustment

14767

Investment in south bay company

14767

(to record parent’s share (100%) of the translation adjustment from the translation of the subsidiary accounts on December 31, 20X1)

Part D

No.

General journal

Debit

Credit

A

Foreign currency transaction loss (7300*(0.64-0.75))

803

Foreign currency units (C$)

803

(to record the exchange loss on foreign currency units hold on December 31, 20X1)